A block of mass 0.50 kg is moving with a speed of 2.00 m/s on a smooth surface. It strikes another mass of 1.00 kg and then they move together as a single body. The energy loss during the collision is

Grade:10

2 Answers

Prudhvi teja
13 years ago

Dear vipin

doing conservation of momentum

new  v = 0.5*2/1+0.5

 = 1/1.5 =2/3

energy loss = k.e final-k.e initila

= 0.5(1.5*(0.66*0.66) - 0.5*2*2)

= -0.6666

therefore energyloss = 0.6666n-m

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly.

All the best.

Regards,
Askiitians Experts
Prudhvi Teja

 

 


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian



Askiitians Expert Bharath-IITD
13 years ago

Dear Vipin,

We know that momentum should be conserved

gives 0.5*2 = (1+0.5)*v

where v i the velocity of entire mass system moving after collision...

v=(1/1.5) m/s

energy lost = kinetic energy of the system before collision - kinetic energy of the system aftr collision

                 = 1/2*0.5*22 -1/2* 1.5 *(1/1.5)2

                 =0.6667 Joules

All the best for IITJEE preparation,

 

With Regards,

Adapa Bharath